# Attention-Aided MMSE for OFDM Channel Estimation: Learning Linear Filters with Attention

### v3

@@ VI-F Performance-Complexity Trade‑off
Fig. 11(b) illustrates the trade-off between the average NMSE and computational complexity across all evaluated TDL channel models. We also summarize the per-inference FLOPs for all the considered methods in Table II. As shown in Fig. 11(b), the proposed A-MMSE achieves superior channel estimation performance with substantially lower computational complexity. Specifically, it yields around $46.3\%$ lower average NMSE and $95.1\%$ fewer FLOPs compared to the average performance of MMSE and Channelformer. The RA-A-MMSE, which incorporates rank adaptation, further reduces complexity with minimal performance loss, consistently outperforming existing approaches.
As discussed earlier, these benefits arise from the fact that A-MMSE relies solely on computationally efficient linear operations during channel estimation. Remarkably, the A-MMSE and the RA-A-MMSE push the boundaries of the performance–complexity trade-off, achieving a new level of estimation accuracy at drastically reduced computational cost. This shift sets a new state-of-the-art for practical channel estimation by redefining the achievable balance between accuracy and efficiency.
TABLE II: FLOPs Comparison for Channel Estimation (${\text{K}}=10^{3},{\text{M}}=10^{6},{\text{G}}=10^{9}$)
@@
#\# of Pilots LL=36 LL=72
MMSE $\approx$11M $\approx$21M
1D-MMSE $\approx$21K $\approx$41K
LS $\approx$7.5K $\approx$15K
ChannelNet $\approx$1.35G $\approx$1.35G
Channelformer $\approx$10M $\approx$21M
@@
VI-G Performance Comparison to Sensing-aided Estimation
Recent sensing-aided channel estimation methods present significant trade-offs compared to the A-MMSE framework, particularly in high-mobility scenarios like TDL-D. The sensing-aided approach requires a complex multi-stage processing pipeline for echo analysis, adding considerable computational overhead. Unlike this, A-MMSE performs channel estimation through a single, efficient linear operation. Moreover, sensing-aided techniques depend critically on echo signal quality, a vulnerability absent in A-MMSE, which maintains consistent performance without auxiliary information. In Fig. 12, we also compare the channel estimation performance between our A-MMSE and the sensing-aided approach. At $15$ dB SNR in TDL-D, A-MMSE achieves an NMSE of $3.26\times 10^{-3}$, outperforming the sensing-aided approach ($5.01\times 10^{-3}$) by $35\%$. This performance gap widens at higher SNRs, where sensing-aided methods plateau while A-MMSE continues to improve. It demonstrates that data-driven learning overcomes the fundamental limitations of echo-based refinement. These results indicate that A-MMSE retains its advantage even without relying on explicit echo information, highlighting its robustness and applicability.
Figure 12: A-MMSE vs. Sensing-aided channel estimation
@@ VII Conclusion
In this paper, we proposed the A-MMSE method for OFDM channel estimation. Our key idea is to integrate the classical SP-based framework with the learning capability of the DNN-based approach. Specifically, the proposed A-MMSE learns the linear filtering matrix via the Attention Transformer and applies it for channel estimation, achieving both high estimation accuracy and substantial computational efficiency. In particular, to enhance the learning of the linear filtering matrix, we developed a novel two-stage Attention encoder that effectively captures the frequency and temporal correlation structures of the OFDM channel. Building on this architecture, we further introduced a rank-adaptive extension of A-MMSE, referred to as the RA-A-MMSE. The RA-A-MMSE stands out by dynamically adjusting the rank of the linear filtering matrix, thereby controlling the computational overhead associated with channel estimation. Through extensive numerical simulations using 3GPP TDL channel models, we demonstrated that the proposed A-MMSE provides significant performance gains in channel estimation compared to conventional SP-based methods (e.g., LS and MMSE) as well as DNN-based methods (e.g., ChannelNet [soltani:commlett:19] and Channelformer [luan2023channelformer]). We further showed that the RA-A-MMSE effectively balances estimation performance and computational complexity, offering substantial reductions in computational overhead with minimal performance degradation.
The proposed A-MMSE and the RA-A-MMSE frameworks offer several key advantages: i) once trained, only the linear filter is used, resulting in architectural simplicity; ii) by leveraging the learning capability of the Attention Transformer, it adapts effectively to diverse channel conditions; and iii) it provides a principled integration of classical signal processing with deep learning. With its modular design and computational efficiency, the A-MMSE aligns well with emerging 6G physical layer design principles. Its integration of data-driven learning with model-based filtering makes it well-suited for standardization in scalable, low-complexity, and AI-native wireless transceiver design. Consequently, the proposed architecture offers a practical pathway toward learning-enhanced, standards-compliant channel estimation in next-generation wireless systems. Future work could explore a promising hybrid model that integrates the A-MMSE framework with sensing-aided approaches, while also developing a practical ground-truth-free training methodology using Ridge regression. These advancements would lay the groundwork for a fully online A-MMSE framework that can adapt in real-time to dynamic wireless environments without perfect channel knowledge.
Appendix A Proof of Kronecker Product Structure
Consider the time-varying channel impulse response $h(t,\tau)$ [Bello1963] with time-frequency transfer function
$L_{H}(t,f)=\int^{\infty}_{-\infty}{h(t,\tau)e^{-j2\pi f\tau}\,d\tau}.$ (38)
Under WSSUS [Bello1963]: $\mathbb{E}\left[h(t,\tau)h^{*}(t^{\prime},\tau^{\prime})\right]=r_{h}(t-t^{\prime};\tau)\delta(\tau-\tau^{\prime}).$ This yields 2D stationarity:
$\mathbb{E}\!\left[L_{H}(t,f)L^{*}_{H}(t^{\prime},f^{\prime})\right]$ (39)
$=\iint\mathbb{E}\!\left[h(t,\tau)h^{*}(t^{\prime},\tau^{\prime})\right]\,e^{-j2\pi f\tau}e^{j2\pi f^{\prime}\tau^{\prime}}\,d\tau d\tau^{\prime}$ (40)
$=\int r_{h}(t-t^{\prime};\tau)\,e^{-j2\pi(f-f^{\prime})\tau}\,d\tau.$ (41)
Setting $\Delta t=t^{\prime}-t$, $\Delta f=f^{\prime}-f$, then we have $\mathbb{E}\left[L_{H}(t,f)L_{H}^{*}(t+\Delta t,f+\Delta f)\right]=R_{H}(\Delta t,\Delta f)$. This correlation function and scattering function form a 2D Fourier pair:
$R_{H}(\Delta t,\Delta f)$ $=\int^{\infty}_{-\infty}{\int^{\infty}_{-\infty}C_{H}(\tau,\nu){e^{j2\pi(\nu\Delta t-\tau\Delta f)\,d\tau d\nu}}}.$ (42)
Assuming separable scattering [Yoojin20152DLMMSE] $C_{H}(\tau,\nu)=P_{\tau}(\tau)\cdot S_{\nu}(\nu)$, we can substitute as
$R_{H}(\Delta t,\Delta f)$ $=\int^{\infty}_{-\infty}{\int^{\infty}_{-\infty}{P_{\tau}(\tau)S_{\nu}(\nu)e^{j2\pi(\nu\Delta t-\tau\Delta f)}}}\,d\tau d\nu$ (43)
$=\int^{\infty}_{-\infty}{P_{\tau}(\tau)e^{-j2\pi\tau\Delta f}\,d\tau}\cdot\int^{\infty}_{-\infty}{S_{\nu}(\nu)e^{j2\pi\nu\Delta t}\,d\nu}$ (44)
$=r_{t}(\Delta t)\cdot r_{f}(\Delta f).$ (45)
For OFDM with $N$ subcarriers and $M$ symbols: $H[n.m]=L_{H}(mT_{sym},\,n\Delta f_{s})$. The discrete correlation: $\mathbb{E}\left[H[n,m]H^{*}[n^{\prime},m^{\prime}]\right]=r_{t}\left[(m-m^{\prime})T_{sym}\right]\cdot r_{f}[(n-n^{\prime})\Delta f_{s}]$ Define frequency and temporal correlation matrices as
  • $\left[\mathbf{R}_{f}\right]_{n,n^{\prime}}=r_{f}\left[(n-n^{\prime})\Delta f_{s}\right]\in\mathbb{C}^{N\times N}$
  • $\left[\mathbf{R}_{t}\right]_{m,m^{\prime}}=r_{t}\left[(m-m^{\prime})T_{sym}\right]\in\mathbb{C}^{M\times M}$
For vectorized channel $\mathbf{h}=\operatorname{vec}\left(\mathbf{H}\right)$ with $h_{n+mN}=H[n,m]$, OFDM channel’s correlation can be represented
$[\mathbf{R}]_{(n+mN),(n^{\prime}+m^{\prime}N)}$ $=\mathbb{E}\left[H[n,m]H^{*}[n^{\prime},m^{\prime}]\right]$ (46)
$=[\mathbf{R}_{t}]_{m,m^{\prime}}\cdot[\mathbf{R}_{f}]_{n,n^{\prime}}.$ (47)
Therefore $\mathbf{R}=\mathbb{E}\left[\mathbf{h}\mathbf{h}^{H}\right]=\mathbf{R}_{t}\otimes\mathbf{R}_{f}$.
Appendix B Performance Penalty of Covariance-Mismatched MMSE
In this appendix, we provide a rigorous mathematical analysis of the performance penalty incurred by covariance-mismatched MMSE estimation. The oracle MMSE filter, assuming perfect knowledge of the true channel covariance $\mathbf{R}$, is given by
$\mathbf{W}^{*}=\Sigma_{xy}\Sigma_{yy}^{-1}=\mathbf{R}\mathbf{S}^{H}\left(\mathbf{S}\mathbf{R}\mathbf{S}^{H}+\sigma^{2}\mathbf{I}\right)^{-1}.$ (48)
where $\mathbf{S}$ denotes the pilot selection matrix, $\Sigma_{xy}=\mathbf{R}\mathbf{S}^{H}$ is the cross-covariance between the full channel and pilot observations, and $\mathbf{S}\mathbf{R}\mathbf{S}^{H}+\sigma^{2}\mathbf{I}$ is the auto-covariance of pilot observations. The corresponding minimum achievable MSE is
$\mathrm{MSE}(\mathbf{W}^{\star})$ $=\mathrm{tr}(\mathbf{R})-\mathrm{tr}\big(\Sigma_{xy}\Sigma_{yy}^{-1}\Sigma_{yx}\big).$ (49)
In practice, we must rely on an estimated or surrogate covariance $\tilde{\mathbf{R}}$, leading to the mismatched filter
$\mathbf{W}_{\mathrm{mis}}=\tilde{\mathbf{R}}\mathbf{S}^{H}\left(\mathbf{S}\tilde{\mathbf{R}}\mathbf{S}^{H}+\sigma^{2}\mathbf{I}\right)^{-1}.$ (50)
The performance of this mismatched filter is strictly worse than the oracle, with total MSE given by
$\mathrm{MSE}(\mathbf{W}_{\mathrm{mis}})=\mathrm{MSE}(\mathbf{W}^{*})+\Delta_{\mathrm{mis}}.$ (51)
where $\Delta_{\mathrm{mis}}\geq 0$ represents the irreducible penalty due to covariance mismatch. The penalty term can be explicitly quantified as
$\Delta_{\mathrm{mis}}=\left\|(\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})\Sigma_{yy}^{1/2}\right\|_{F}^{2}\geq 0.$ (52)
To prove this, we expand the MSE of the mismatched filter:
$\mathrm{MSE}(\mathbf{W}_{\mathrm{mis}})=\mathbb{E}\!\left[\left\|\mathbf{H}-\mathbf{W}_{\mathrm{mis}}\mathbf{y}_{p}\right\|_{2}^{2}\right].$ (53)
where $\mathbf{H}$ is the true channel vector and $\mathbf{y}_{p}$ is the pilot observation vector. By adding and subtracting $\mathbf{W}^{\star}\mathbf{y}_{p}$, we obtain
$\mathrm{MSE}(\mathbf{W}_{\mathrm{mis}})=\mathbb{E}\!\left[\left\|(\mathbf{H}-\mathbf{W}^{*}\mathbf{y}_{p})+(\mathbf{W}^{*}-\mathbf{W}_{\mathrm{mis}})\mathbf{y}_{p}\right\|_{2}^{2}\right].$ (54)
Using the orthogonality principle of the optimal MMSE estimator, the cross term vanishes, yielding
$\mathrm{MSE}(\mathbf{W}_{\mathrm{mis}})=\mathrm{MSE}(\mathbf{W}^{*})+\mathbb{E}\!\left[\left\|(\mathbf{W}^{*}-\mathbf{W}_{\mathrm{mis}})\mathbf{y}_{p}\right\|_{2}^{2}\right].$ (55)
The second term can be written as
$\Delta_{\mathrm{mis}}$ $=\operatorname{tr}\!\left((\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})\boldsymbol{\Sigma}_{yy}(\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})^{\sf{H}}\right).$ (56)
which is equivalent to $\Delta_{\mathrm{mis}}=\left\|(\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})\Sigma_{yy}^{1/2}\right\|_{F}^{2}\geq 0.$
For small covariance mismatches $\Delta\mathbf{R}=\tilde{\mathbf{R}}-\mathbf{R}$, we can establish a local quadratic lower bound. Let $\tilde{{\Sigma}}_{yy}=\mathbf{S}\tilde{\mathbf{R}}\mathbf{S}^{\sf{H}}+\sigma^{2}\mathbf{I}.$ and $\Delta{\Sigma}_{yy}=\tilde{{\Sigma}}_{yy}-{\Sigma}_{yy}=S\,\Delta\mathbf{R}\,\mathbf{S}^{\sf{H}}.$. Using first-order perturbation analysis of the matrix inverse, we have
$\mathbf{W}_{\text{mis}}-\mathbf{W}^{*}\approx\Delta\mathbf{R}\mathbf{S}^{\sf{H}}{\Sigma}_{yy}^{-1}-\mathbf{R}\mathbf{S}^{\sf{H}}{\Sigma}_{yy}^{-1}\Delta{\Sigma}_{yy}{\Sigma}_{yy}^{-1}.$ (57)
Substituting $\Delta{\Sigma}_{yy}=\mathbf{S}\,\Delta\mathbf{R}\,\mathbf{S}^{\sf{H}}$ and simplifying as
$\mathbf{W}_{\text{mis}}-\mathbf{W}^{*}\approx(\mathbf{I}-\mathbf{W}^{*}{\mathbf{S}})\,\Delta\mathbf{R}\,\mathbf{S}^{\sf{H}}{\Sigma}_{yy}^{-1}.$ (58)
Therefore, the mismatch penalty satisfies
$\Delta_{\text{mis}}\geq\lambda_{\min}({\Sigma}_{yy})\left\|(\mathbf{I}-\mathbf{W}^{*}\mathbf{S})\,\Delta\mathbf{R}\,\mathbf{S}^{\sf{H}}{\Sigma}_{yy}^{-1/2}\right\|_{F}^{2}.$ (59)
Using the matrix norm property $\lVert\mathbf{AB}\rVert_{F}\geq\sigma_{\mathrm{min}}(\mathbf{A})\lVert\mathbf{B}\rVert_{F}$, where $\sigma_{\mathrm{min}}$ denotes the smallest singular value of a matrix, we obtain
$\Delta_{\text{mis}}\geq\lambda_{\min}({\Sigma}_{yy})\,\sigma_{\min}^{2}(\mathbf{I}-\mathbf{W}^{*}\mathbf{S})\,\sigma_{\min}^{2}(\mathbf{S})\left\|{\Sigma}_{yy}^{-1/2}\Delta\mathbf{R}\right\|_{F}^{2}.$ (60)
Further, noting that $\left\|{\Sigma}_{yy}^{-1/2}\Delta\mathbf{R}\right\|_{F}^{2}\geq\frac{\left\|\Delta\mathbf{R}\right\|_{F}^{2}}{\rho_{\mathbf{S}}^{2}\left\|\mathbf{R}\right\|_{2}+\sigma^{2}}$ for $\mathbf{S}$ with a bounded spectral norm $\rho_{s}=\lVert\mathbf{S}\rVert_{2}$, we establish the quadratic lower bound
$\Delta_{\text{mis}}\geq\lambda_{\min}({\Sigma}_{yy})\,c^{2}\left\|\Delta\mathbf{R}\right\|_{F}^{2},$ (61)
where $c=\frac{\sigma_{\min}(\mathbf{I}-\mathbf{W}^{*}\mathbf{S})\,\sigma_{\min}(\mathbf{S})}{\rho_{s}^{2}\left\|\mathbf{R}\right\|_{2}+\sigma^{2}}>0$ is a positive constant dependent on the system parameters. Even small covariance estimation errors result in an irreducible quadratic performance penalty for conventional signal processing methods.

### v4

@@ VI-D Performance-Complexity Trade‑off
Fig. 8 illustrates the trade-off between the average NMSE and computational complexity, averaged across all SNR levels and both HSR and SU scenarios. We also summarize the per-inference FLOPs for all the considered methods in Table II. As shown in Fig. 8, the proposed A-MMSE achieves superior channel estimation performance with substantially lower computational complexity. Specifically, it yields around $64.5\%$ lower average NMSE and $95.1\%$ fewer FLOPs compared to the average performance of LMMSE and Channelformer.
The rank-adaptive extension further optimizes this efficiency. The RA-A-MMSE with $10\%$ rank retains high accuracy despite the reduced rank, achieving approximately $50.5\%$ lower average NMSE compared to Channelformer. This demonstrates that the RA-A-MMSE effectively exploits the estimator’s low-rank structure, offering a scalable solution that outperforms DNN baselines with reduced computational overhead. As discussed, these benefits arise from the fact that A-MMSE relies solely on computationally efficient linear operations during channel estimation. Consequently, the A-MMSE and RA-A-MMSE frameworks advance the performance–complexity trade-off, achieving improved estimation accuracy at significantly reduced computational cost. This establishes a practical benchmark for channel estimation by enhancing the trade-off between accuracy and efficiency.
TABLE II: FLOPs Comparison for Channel Estimation ($L=72$)
@@
LMMSE $\approx 21$ M
1D-LMMSE $\approx 41$ K
LS $\approx 15$ K
ChannelNet $\approx 1.35$ G
Channelformer $\approx 21$ M
@@ VII Conclusion
In this paper, we proposed the A-MMSE method for OFDM channel estimation. Our key idea is to integrate the classical SP-based framework with the learning capability of the DNN-based approach. Specifically, the proposed A-MMSE learns the linear filtering matrix via the Attention Transformer and applies it for channel estimation, achieving both high estimation accuracy and substantial computational efficiency. In particular, to enhance the learning of the linear filtering matrix, we developed a novel two-stage Attention encoder that effectively captures the frequency and temporal correlation structures of the OFDM channel. Building on this architecture, we further introduced a rank-adaptive extension of A-MMSE, referred to as the RA-A-MMSE. The RA-A-MMSE is distinguished by dynamically adjusting the rank of the linear filtering matrix, thereby controlling the computational overhead associated with channel estimation. Through extensive numerical simulations using COST2100 channel models, we demonstrated that the proposed A-MMSE provides significant performance gains in channel estimation compared to conventional SP-based methods as well as DNN-based methods. We further showed that the RA-A-MMSE effectively balances estimation performance and computational complexity, offering substantial reductions in computational overhead with minimal performance degradation.
The proposed A-MMSE and the RA-A-MMSE frameworks offer several key advantages: i) once trained, only the linear filter is used, resulting in architectural simplicity; ii) by leveraging the representation learning capability of the Attention Transformer, they effectively captures the intrinsic correlation structure of wireless channels; and iii) they provide a principled integration of classical signal processing with deep learning. With its efficient design, the A-MMSE aligns well with emerging 6G physical-layer principles, emphasizing scalability and AI-native architectures.
Future work may extend this framework to scenarios where ground-truth channel information is unavailable during training. In addition, exploring reinforcement learning paradigms that leverage end-to-end performance metrics, such as log-likelihood ratios (LLRs), represents a promising direction. Extending our framework to MIMO-OFDM channel acquisition [16] is also interesting.